OpenStudy (anonymous):

Train a and b are traveling in the same direction on parallel tracks train a is traveling in the direction on parallel tracks . train a is traveling at 80 miles per hour b is traveling at 84 miles per hour . Train a passes a station at 9:25 a.m if train b passes the same station at 9:55 am what time will train b catch up with train a ?

OpenStudy (anonymous):

speed of train A = 80 mph speed of train B = 84 mph since train A leaves 30 min earlier and is having speed = 80 mph, it has already covered a distance of 40 miles before train B starts from there let the two trains meet after time = x h distance covered by train A in x h = 80x + 40 distance covered by train B in x h = 84x catching up means that both have covered same distance => 84x = 80x + 40 => 84x - 80x = 40 => 4x = 40 => x = 40/4 = 10 So train B will catch up train A after 10 hours i.e 10 hours after 9.55 am i.e. at 7:55 pm